# Magneto-optical trap loading with an effusive oven in    a large optical access experiment

### v1

@@
Magneto-optical trap loading with an effusive oven in    a large optical access experiment
M. Gaudesius1,2, [^1] Contact author: marius.gaudesius-1@ou.edu
@@
We present an experimental, numerical, and analytical study of a strontium magneto-optical trap (MOT) loaded from an effusive oven in a configuration optimized for high numerical aperture optical tweezers. Our approach orients the cold atom flux along the MOT symmetry axis to reduce the experimental complexity and maximize the overall optical access into the scientific region of study. We use a moving molasses technique to enable this configuration and show that its performance depends critically on metastable-state shelving (to $5s5p\,^{3}\!P_{2}$) during the atom transfer to the 3D MOT. Furthermore, we find that the parameters for optimal transfer efficiency are bounded by dark-state loss (to $5s5p\,^{3}\!P_{0}$) in the trap region where repumping is present. These observations are verified to great degree of accuracy using both our developed analytical and numerical models. The corresponding 3D simulation tool is used to perform a comprehensive study of the trap loading dynamics, beginning at the oven exit and ending at the 3D MOT, demonstrating its effectiveness in optimizing an effusive oven experiment.
@@ I. Introduction
Both alkali and alkaline-earth metal atoms form an important part of cold atom experiments owing to easily accessible transition frequencies and effectively closed cooling cycles. The latter atoms, possessing two valance electrons instead of one, have unique energy level structures that have attracted a growing interest in the atomic physics community. For example, a bosonic isotope of strontium, 88Sr, offers a pure ground state and multiple narrow transitions, making it an outstanding candidate in several areas of research, such as atom interferometry [3, 4], optical frequency standards [5, 6], tests of fundamental physics [7, 8], as well as quantum simulation and computation with cold Rydberg atoms [9, 10, 11]. These kinds of investigations generally are based on techniques that are widely used and established. Particularly for the Rydberg-atom based simulation and computing, the manipulation and control of atoms relies heavily on the usage of optical tweezers [12]-[15]. In such experiments, it is often desirable to achieve a sizable atomic array [16, 17], a high detection efficiency via large solid-angle light-collection [18, 19], and an extensive trap-lifetime [20, 21].
@@
Modern experiments commonly accomplish this by using high numerical aperture (NA) microscope objectives to create and detect a tweezer array, with the atoms loaded from a magneto-optical trap (MOT) in UHV conditions. The high NA objectives consume precious optical access, often leading to a departure from the optimal MOT laser beam configuration [22, 23]. Loading a MOT from a cold flux source achieves UHV conditions and increased experimental repetition rates [24, 25, 26], but fur-
ther consumes lines of access into the scientific study region of the apparatus. We propose and demonstrate in this paper a technique for the MOT loading that solves this challenge in an effusive oven experiment. Hence, high NA optical tweezers can be realized while simultaneously opening additional laser-pathways for atom array manipulation. We verify and optimize this  approach  to  great accuracy with a simulation tool incorporating the physics deemed by our analytical theory to be critical for our main observations.
@@
A variety of methods have been demonstrated for the MOT loading [27]-[31].  For  high  vapor  pressure  atoms such as rubidium or cesium, a MOT can be loaded directly from background vapor [32, 33]. On the other hand, for low vapor pressure atoms such as strontium or ytterbium, an effusive oven [34, 35], or other methods [36, 37], must be used to achieve sufficient density. The effusive oven approach typically relies on collimating the atoms into a cold jet that is directed at a MOT. For reduced collisions, a two-stage MOT loading method can be taken, where the atoms from the original MOT (commonly, 2D) are transferred to the main MOT (3D) with the help of a push beam. The transfer can be done through a differential pumping stage, leading to an improved vacuum in the science region [38, 39] and hence achieving a long trap lifetime required for extended manipulation and detection of large atom arrays [18, 40, 41].
@@
The effusive oven approach is not without its challenges. Heating the reservoir to a large temperature exceeding, e.g., $400\;^{\circ}$ for strontium, requires regular refurbishments of the chamber components as well as a height-
@@
ened maintenance of the vacuum [39, 41, 42]. Moreover, a Zeeman slower [43, 44] is required to collect a large fraction of the hot atoms to within the capture speed of a downstream MOT. In a two-stage MOT method, the loading efficiency of a 3D MOT is bounded by that of a 2D MOT and is critically dependent on a push (transfer) beam [38, 39, 45, 46]. A growing list of  lines of  access into the science region is thus obtained, which can  include (i) cold flux/push beam, (ii) six 3D MOT beams, (iii) one or two high NA microscope objectives, and (iv) one or more lasers for manipulating the quantum state of the array. In our approach, this challenge is ameliorated by creating the cold flux with two of the 3D MOT beams, hence removing the push beam from the list. The optimization of this approach requires detailed considerations of an asymmetric MOT configuration with complex loading dynamics involving metastable-state shelving and dark-state loss. Hence, a pivoting goal of this paper is to provide a comprehensive simulation tool addressing the dynamics associated with the loading of MOTs, beginning at the hot source and ending at the 3D MOT. An analytical theory that we develop provides a clear-cut physical picture behind our 3D MOT loading approach.
We start in Sec. II by giving an overview of our effusive strontium-oven experimental setup that is optimized for high NA optical tweezers and discuss the corresponding numerical setup. It is based  on  a  model  accounting
@@
for light polarization, arbitrary beam orientations in 3D space, and atom lifetime. In Sec. III, we describe our technique for populating the main science region while excluding an additional push beam, and present the main measurement results with comparison to our numerical and analytical models. Sec. IV presents a supplementary numerical study of the loading, including general guidelines for optimizing an effusive oven experiment. In Sec. V, we conclude and discuss the future prospects of our work. Apps. A1 and A2 cover the theoretical aspects of our numerical model, and App. B provides analytical calculations in support of our main observations.
@@ II. Experimental and numerical
Our experiment employs the bosonic isotope 88Sr (mass $m=1.46\cdot 10^{-25}$ kg), operating on the blue transition $5s^{2}\,{}^{1}\!S_{0}\rightarrow 5s5p\,^{1}\!P_{1}$ with the wavelength $\lambda=461$ nm, the linewidth $\Gamma=2\pi\times 32$ MHz, and the saturation intensity $I_{sat}=42.72$ mW/cm2. The effusive atomic beam cooling, and the magneto-optical trapping together with the moving molasses transfer are thus realized with radiation pressure from this ${}^{1}\!S_{0}\;{\rightarrow}\;^{1}\!P_{1}$ transition, as discussed in detail below. The atoms exposed to the blue light can shelve to the metastable stable $5s5p\,^{3}\!P_{2}$ (reached via  the
[…]
@@
The experimental setup is seen in Fig. 1(a). Its main stages include an effusive oven, a 2D optical molasses (OM), a zeeman slower (ZS), a 2D MOT, and, lastly, a 3D MOT. The atoms originating from the effusive oven are first collimated to form a jet that is cooled transversely by the 2D OM and longitudinally by the ZS. The longitudinal cooling slows the atoms into a velocity class within the capture range of the 2D MOT, with the transverse cooling providing additional collimation for an improved loading efficiency. The 3D MOT is loaded from the 2D MOT by using the axial beams only (discussed in Sec. III), granting the optical access for the high NA tweezer objectives. We use a 240 l/s ion-getter pump (SAES NEXTorr Z 200) that we attach above the 2D MOT chamber. Without the getter activated, it maintains an estimated $10^{-9}$ Torr pressure at the main science region. With this pressure, the mean lifetime due to background collisions is expected to be on the order of 1 s [51], which we deem satisfactory for our investigations. We determine the atom numbers at various chamber locations (2D OM, 2D MOT, and 3D MOT) from the voltage readings of an avalanche photodiode (Hamamatsu C12703-01) together with calculated scattering rates and detection solid angles [52, 53].
@@
The numerical setup is seen in  Fig.  1(b).  It  mimics the experimental one starting from the 2D OM stage, with its initial numerical conditions dictated by the collimation geometry after the oven exit and the measured flowrate (Sec. II.A). We base our model on the $F=0\rightarrow F^{\prime}=1$ transition, faithful to blue trapping descriptions of 88Sr and allowing for a proper treatment of the features related to the magnetic field and light polarization. We build upon the model in Ref. [54] by considering arbitrary beam orientations, although we currently exclude its multiple-scattering effects. The theoretical aspects are detailed in Apps. A1 and A2, where the radiation pressure and the atom lifetime are respectively treated.  The
[…]
@@ A. Strontium oven and
We operate the strontium oven at a relatively modest temperature of $440\,^{\circ}$, in an effort to minimize the chamber degradation over time. We stack microcapillaries (at $500\,^{\circ}$, higher to prevent clogging) collimating the atoms into a jet [57] and use a differential tube (before the 2D OM) providing additional collimation that results in the maximal divergence angle $\theta_{max}=23$ mrad at an opening of 6 mm and the flowrate of $2\times 10^{11}$ atoms/s measured at the 2D OM location. The 2D OM itself is realized by two retroreflected beams with an individual power of 6 mW, a $1/e^{2}$ radius of $3.2$ mm, and a detuning of $-30$ MHz.
The oven temperature, the maximal divergence angle, and the atomic flowrate set the initial conditions for  the
2D OM simulation. The flowrate is our only free numerical parameter, due to it being entirely dependent on a measurement. Each superparticle in the flow has a corresponding divergence angle found by drawing from a normal distribution with $\mu_{f}=2\sigma_{f}=\theta_{max}/2$, where $\mu_{f}$ and $\sigma_{f}$ are the distribution mean and standard deviation, respectively; if the result is smaller than 0 or larger than $\theta_{max}$, a uniform distribution draw is made between these limits. This choice is based on a geometrical reasoning that the atoms from a given capillary have an equal probability to emerge at the divergence angle of $\theta_{max}/2$, while for those away from the center there exists a decreasing probability to emerge at greater angles; the angles smaller than $\theta_{max}/2$ are treated the same way as the larger ones for simplicity. A calculated divergence angle finally determines a superparticle’s initial transverse velocity from the total velocity (sum of longitudinal and transverse) that is drawn from the Maxwell-Boltzmann distribution evaluated at the oven temperature. Note that, in general, by introducing a geometry, the velocity distribution becomes modified. To first approximation, we expect our choice to hold true. The effect of the 2D OM is evaluated at the downstream 2D MOT, as discussed in Sec. II.C, while the numerical optimization is discussed in Sec. IV.

### v2

# Magneto-optical-trap loading in a large optical-access experiment

@@
Magneto-optical-trap loading in a large optical-access experiment
[…]
  Contact author: gd.mar900@gmail.com
@@
We present an experimental, numerical, and analytical study of strontium magneto-optical trap (MOT) loading from a cold atomic beam in a configuration optimized for high numerical aperture optical tweezers. Our approach orients the beam flow along the MOT symmetry axis to reduce the experimental complexity and maximize the overall optical access into the scientific region of study. We use a moving molasses technique to enable this configuration and show that its performance depends critically on metastable-state shelving (to $5s5p\,^{3}\!P_{2}$) during the atom transfer to the three-dimensional (3D) MOT. Furthermore, we find that the parameters for optimal transfer efficiency are bounded by dark-state loss (to $5s5p\,^{3}\!P_{0}$) in the trap region where repumping is present. These observations are verified to great degree of accuracy using both our developed analytical and numerical models. The corresponding 3D simulation tool is used to perform a comprehensive study of the trap loading dynamics, beginning at the oven exit and ending at the 3D MOT, demonstrating its effectiveness in optimizing an effusive oven experiment.
@@ I. Introduction
Both alkali and alkaline-earth metal atoms form an important part of cold atom experiments owing to easily accessible transition frequencies and effectively closed cooling cycles. The latter atoms, possessing two valance electrons instead of one, have unique energy level structures that have attracted a growing interest in the atomic physics community. For example, a bosonic isotope of strontium, 88Sr, offers a pure ground state and multiple narrow transitions, making it an outstanding candidate in several areas of research, such as atom interferometry [3, 4], optical frequency standards [5, 6], tests of fundamental physics [7, 8], as well as quantum simulation and computation with cold Rydberg atoms [9, 10, 11]. These kinds of investigations generally are based on techniques that are widely used and established. Particularly for the Rydberg-atom based simulation and computing, the manipulation and control of atoms relies heavily on the usage of optical tweezers [12]-[15]. In such experiments, it is often desirable to achieve a sizable atomic array [16, 17], a high detection efficiency via large solid-angle light collection [18, 19], and an extensive trap lifetime [20, 21].
@@
Modern experiments commonly accomplish this by using high numerical aperture (NA) microscope objectives to create and detect a tweezer array, with the atoms loaded from a magneto-optical trap (MOT) in UHV conditions. The high NA objectives consume precious optical access, often leading to a departure from the optimal MOT laser beam configuration [22, 23]. Loading a MOT from a cold source achieves UHV conditions and increased experimental repetition rates [24, 25, 26], but further consumes lines of access into the scientific study region of the apparatus. We propose and demonstrate in this paper a technique for the MOT loading that solves this challenge in an effusive oven experiment. Hence, high NA optical tweezers can be realized while simultaneously opening additional laser pathways for atom array manipulation. We
verify and optimize this approach to great accuracy with a simulation tool incorporating the physics deemed by our analytical theory to be critical for our main observations.
Figure 1: Moving molasses technique for transferring atoms between magneto-optical traps (MOTs) and enabling large optical access experiments (with proposed locations for microscope objectives). The 3D MOT loading relies on its axial beams creating an atomic molasses beam from the location of the 2D MOT (itself loaded by, e.g., a Zeeman-slowed Sr flux). We present two cases: (i) Aligned, where the top and bottom axial beams are counter propagating and must be unequally detuned ($\delta_{\text{t}}\neq\delta_{\text{b}}$); and (ii) Misaligned, where the bottom beam is slightly angled (to 3D MOT) for enhanced loading with $\delta_{\text{t}}\neq\delta_{\text{b}}$, in addition to opening the possibility to load given $\delta_{\text{t}}=\delta_{\text{b}}$. Note that the results of this work consider case (i) unless specified.
@@
A variety of methods have been demonstrated for the MOT loading [27]-[31]. For high vapor pressure atoms such as rubidium or cesium, a MOT can be loaded directly from background vapor [32, 33]. On the other hand, for low vapor pressure atoms such as strontium or ytterbium, an effusive oven [34, 35] or other methods [36, 37] must be used to achieve sufficient density. The effusive oven approach typically relies on collimating the atoms into a beam that is directed at a MOT. For reduced collisions, a two-stage MOT loading method can be taken, where the atoms from the original MOT (commonly, 2D) are transferred to the main MOT (3D) with the help of different techniques (see below). The transfer can be done through a differential pumping stage, leading to an improved vacuum in the science region [38, 39] and hence achieving a long trap lifetime required for the extended manipulation and detection of large atom arrays [16, 40, 41].
@@
The effusive oven approach is not without its challenges. Heating the reservoir to a large temperature exceeding, e.g., $400\;^{\circ}$C for strontium requires regular refurbishments of the chamber components as well as heightened maintenance of the vacuum [39, 41, 42]. Moreover, a Zeeman slower [43, 44] is required to collect a large fraction of the hot atoms to within the capture speed of a downstream MOT. In a two-stage MOT method, the loading efficiency of a 3D MOT is bounded by that of a 2D MOT and is critically dependent on the transfer technique that is employed.
A simple way of performing the transfer is by adding to the 2D MOT a single push beam that creates an atomic flow towards the 3D MOT [38, 39, 45, 46]. For enhanced flux, a 2D+ MOT can be used, where a pair of Doppler-cooling “+” beams are added orthogonally to the 2D MOT beams, and an atomic beam is created within a shadow of the “+” beam that opposes the pushing “+” beam [24]. This shadow or extraction column is produced by a small aperture in the in-vacuum mirror that reflects light to create the opposing beam. This approach is similar to the low-velocity intense source (LVIS) system [47], where the atomic beam emerges from an aperture in a MOT mirror. The transfer technique employing a 2D+ MOT achieves the flow with the “+” beams having different intensities, while the LVIS system uses the same beam parameters but results in weaker extraction. In contrast to these approaches, we utilize differently detuned 3D MOT axial beams and no in-vacuum assistance (see Fig. 1), allowing for relatively straightforward assembly of our system and modeling of the transfer from our 2D MOT. We note that it is not called a 2D+ MOT to stress that the axial beams are part of the 3D MOT system and that the axial beam misalignment is a viable option for transfer (misaligned case in Fig. 1). In the past [48], the
misalignment of 2D MOT beams themselves has been used to achieve the transfer into the 3D MOT region but without improved optical access to this main region like in the case of all the previous techniques. Overall, our approach creates broad optical access to the 3D MOT that uses convenient beam geometry achieved by having its axial beams creating the flow from the 2D MOT while excluding in-vacuum parts.
@@
The optimization of our approach requires detailed considerations of an asymmetric MOT configuration with complex loading dynamics involving metastable-state shelving and dark-state loss. Hence, a pivotal goal of this paper is to provide a comprehensive simulation tool addressing the dynamics associated with the loading of MOTs, beginning at the hot source and ending at the 3D MOT. An analytical theory that we develop provides a clear-cut physical picture behind our 3D MOT loading approach.
@@
We start in Sec. II by giving an overview of our effusive strontium-oven experimental setup that is optimized for high NA optical tweezers and discuss the corresponding numerical setup. It is based on a model accounting for light polarization, arbitrary beam orientations in 3D space, and atom lifetime. In Sec. III, we describe our technique for populating the main science region and present the main measurement results with comparison to our numerical and analytical models. In Sec. IV, we conclude and discuss the future prospects of our work. Appendixes A and B cover theoretical aspects of our numerical model, and Appendix C provides analytical calculations in support of our main observations. Additionally, the Supplemental Material [49] includes a numerical study of the loading, with general guidelines for optimizing an effusive oven experiment.
@@ II. Experimental and numerical
Our experiment employs the bosonic isotope 88Sr (mass $m=1.46\cdot 10^{-25}$ kg), operating on the blue transition $5s^{2}\,{}^{1}\!S_{0}\rightarrow 5s5p\,^{1}\!P_{1}$ with the wavelength $\lambda=461$ nm, the linewidth $\Gamma=2\pi\times 32$ MHz, and the saturation intensity $I_{sat}=42.72$ mW/cm2. The effusive atomic beam cooling, and the magneto-optical trapping together with the moving molasses transfer are thus realized with radiation pressure from this $^{1}\!S_{0}\;{\rightarrow}\;^{1}\!P_{1}$ transition, as discussed in detail below. The atoms exposed to the blue light can shelve to the metastable stable $5s5p\,^{3}\!P_{2}$ (reached via the decay channel $5s5p\,^{1}\!P_{1}\rightarrow 5s4d\,^{1}\!D_{2}\rightarrow\,5s5p$ $^{3}\!P_{2}$), which we find to be critical for achieving an efficient transfer into the final trapping location (from the  2D  MOT  into
the 3D MOT). Repumping is employed only at the 3D MOT and tuned to the cyan transition $5s5p\,^{3}\!P_{2}\rightarrow 5p^{2}\,{}^{3}\!P_{2}$ with the wavelength $\lambda_{c}=481$ nm, the linewidth $\Gamma_{c}=2\pi\times 14$ MHz, and the saturation intensity $I_{sat,c}=16.41$ mW/cm2. This leads to the shelved atom revival, increasing the 3D MOT lifetime [38, 50, 51] and enhancing the detection fidelity in optical tweezer experiments [23, 52]. Note that a loss into the long-lived dark state $5s5p\,^{3}\!P_{0}$ can occur during the repumping, bounding the lifetime increase [50, 51].
@@
The experimental setup is seen in Fig. 2(a). Its main stages include an effusive oven, a 2D optical molasses (OM), a Zeeman slower (ZS), a 2D MOT, and, lastly, a 3D MOT. The atoms originating from the effusive oven are first collimated to form a beam that is cooled transversely by the 2D OM and longitudinally by the ZS. The longitudinal cooling slows the atoms into a velocity class within the capture range of the 2D MOT, with the transverse cooling providing additional collimation for improved loading efficiency. The 3D MOT is loaded from the 2D MOT by using the axial beams only (discussed in Sec. III), granting the optical access for the high NA tweezer objectives. We use a 240 l/s ion-getter pump (SAES NEXTorr Z 200) that we attach above the 2D MOT chamber. Without the getter activated, it maintains an estimated $10^{-9}$ Torr pressure at the main science region. With this pressure, the mean lifetime due to background collisions is expected to be on the order on 1 s [53], which we deem satisfactory for our investigations. We determine the atom numbers at various chamber locations (2D OM, 2D MOT, and 3D MOT) from the voltage readings of an avalanche photodiode (Hamamatsu C12703-01)
together with calculated scattering rates and detection solid angles [34, 54].
@@
The numerical setup is seen in Fig. 2(b). It mimics the experimental one starting from the 2D OM stage, with its initial numerical conditions dictated by the collimation geometry after the oven exit and the measured flow rate (Sec. II.A). We base our model on the $F=0\rightarrow F^{\prime}=1$ transition, faithful to the blue trapping descriptions of 88Sr and allowing for a proper treatment of the features related to the magnetic field and light polarization. We build upon the model in Ref. [55] by considering arbitrary beam orientations, although we currently exclude its multiple-scattering effects. The theoretical aspects are detailed in Appendixes A and B, where the radiation pressure and the atom lifetime are respectively treated.
The dynamics are numerically implemented using the Leapfrog algorithm that updates the velocities and positions of the superparticles (collections of regular particles) once the forces are computed (scaled by the number of regular particles represented by one superparticle) [55, 56]. To conserve computational resources, the simulation is divided into three parts (with the first two determining the initial conditions for their respective next part): (i) 2D OM, (ii) ZS and 2D MOT, and (iii) 2D and 3D MOTs. [Figure 2(b) showcases the simulation after (i), with (ii) and (iii) performed simultaneously.] We use up to $10^{7}$ superparticles, in that way obtaining convergent results, and release them continuously according to the given loading rates (at the simulation end, all the superparticles have been released). Each superparticle has a lifetime calculated by multiplying the mean lifetime (Eq. B1) with a number predrawn from a unit exponential dis-
tribution. Once the lifetime is over in a region without the repumping (the enhancement factor $\mathcal{E}=1$ in Eq. B1), a superparticle is affected only by gravity ($\textsf{z}$ direction), in accordance with it being shelved to the metastable state $5s5p\,^{3}\!P_{2}$. Given a superparticle drifts into the repumping region, it is revived with a new lifetime (with $\mathcal{E}=27$) after which it is deleted from the simulation as it is assumed to be lost to the dark state $5s5p\,^{3}\!P_{0}$. Note that this new lifetime assumes an enhancement ($\mathcal{E}$) that is approximately maximal for our employed repumping scheme and is independent of the atomic density (which changes during MOT loading) [51]. Note also that the superparticles are deleted at the chamber bounds in accordance with the surface binding energy of strontium being especially large [57], and the beams naturally have a finite diameter as set by the size of the optics apertures (half-inch). In the below subsections, we discuss in detail the experimental and numerical setups.
@@ A. Strontium oven and
We operate the strontium oven at a relatively modest temperature of $440\,^{\circ}$C, in an effort to minimize the chamber degradation over time. We stack microcapillaries (at $500\,^{\circ}$C, higher to prevent clogging) collimating the atoms into a beam [58] and use a differential tube (before the 2D OM) providing additional collimation. Each microcapillary is $l_{cap}=8$ mm long with the inner diameter of $d_{cap}=0.17$ mm, while the differential tube is $l_{dif}=44$ mm long with the inner diameter of $d_{dif}=6$ mm; the separation distance between the microcapillary stack and the differential tube is $\Delta l=204$ mm. These dimensions of our dual stage collimation affect the transverse velocity profile as discussed below. The flow rate measured at the location of the 2D OM is $2\times 10^{11}$ atoms/s. The 2D OM itself is realized by two retroreflected beams with an individual power of 6 mW, a $1/e^{2}$ radius of $3.2$ mm, and a detuning of 30-30 MHz.
The initial conditions for the 2D OM simulation are set by (i) the oven temperature; (ii) the atomic flow rate; and (iii) the collimation divergence angle calculated using the dual stage collimation length $l_{tot}\equiv l_{cap}+l_{dif}+\Delta l$ and the end opening size $d_{dif}$, $\theta_{d}=\text{tan}\left(\frac{d_{dif}}{l_{tot}}\right)\approx 23$ mrad (which by construction lies between the maximum divergence angles from the capillary start and end, respectively, $\text{tan}\left(\frac{d_{cap}}{l_{cap}}\right)\approx 21$ mrad and $\text{tan}\left(\frac{d_{dif}}{l_{tot}-l_{cap}}\right)\approx 24$ mrad). The flow rate is our only free numerical parameter, because it is entirely dependent on a measurement. The divergence-angle distribution for the superparticles is a combination of (i) a normal distribution with $\mu_{d}=2\sigma_{d}=\theta_{d}/2$ ($\mu_{d}$ and $\sigma_{d}$ are the mean and standard deviation) and cut off outside 0 and $\theta_{d}$, and (ii) a uniform distribution that lies between these cut off values and that is relevant only for the superparticles where the initial normal distribution draw exceeds these cut offs. A given divergence angle from the distribution finally  deter-
mines a superparticle’s initial transverse velocity from the total velocity (sum of longitudinal and transverse) that is drawn from the Maxwell-Boltzmann distribution evaluated at the oven temperature. The effect of the 2D OM is evaluated at the 2D MOT, as discussed in Sec. II.C, while the numerical optimization is discussed in the Supplemental Material [49].
Figure 3: Numerically calculated radial velocity $v_{r}$ distributions after our dual stage collimation for strontium particles emerging from an oven. The dark gray histogram is obtained using our divergence-angle assumption, whereas the light gray histogram is obtained with MolFlow+ [59]. See the main text for details.
To justify the choice of our divergence-angle distribution, we use MolFlow+ [59] to perform a simulation of oven-particle trajectories. This simulation is initiated inside of a cube (20 mm edge length) representing an oven, and the trajectories are let to evolve through our dual stage collimation geometry until a total of $10^{4}$ events are recorded at the output. The respective temperatures of the oven and the capillaries are as stated above ($440\,^{\circ}$C and $500\,^{\circ}$C), while the differential tube is at room temperature (as in the experiment). The initial velocities are drawn from the Maxwell-Boltzmann distribution, and the angles of reflection by surfaces are drawn from the cosine distribution (according to Knudsen’s cosine law). No sticking to surfaces is assumed, but hitting the tube results in practically infinite dwell-time due to the much smaller tube temperature. This effectively filters away the trajectories with divergence angles larger than those resulting from reflections right off the periphery of a given capilary and continuing near the tube end edge. We note that the corresponding greatest angle is roughly equal to our own model’s upper cutoff (comparing ${\approx}24$ to ${\approx}23$ mrad). In Fig. 3, we compare the distributions for the output radial velocities $v_{r}$ obtained using our divergence-angle assumption (dark gray) and MolFlow+ (light gray). We observe that the two distributions are nearly identical (up to randomization error), proving that our simple assumption can accurately predict the physical velocities for our geometry output.